# Should We Still Pretrain Encoders with Masked Language Modeling?

### v2

@@ 4 Two-Stage CLM+MLM Pretraining
Building on Section 3, which shows that CLM outperforms MLM in TC performance, data efficiency, and training stability, we explore the benefits of applying CLM and MLM sequentially during pretraining.
Under fixed compute constraints, starting pretraining with CLM and continuing with MLM yields better results than MLM alone.
Motivated by the strong performance of CLM in learning text representations, we investigate the impact of two-stage pretraining objectives, starting with CLM and transitioning to MLM. We evaluate this under fixed compute budgets of 12,000, 22,000, and 42,000 training steps, using different splits: 100% MLM, 25%-75%, 50%-50%, 75%-25%, and 100% CLM. Interestingly, as shown in Figure 6, combining CLM and MLM consistently improves downstream performance compared to using MLM alone, though the effect varies by task and training budget. Overall, a split between 25%-75% and 50%-50% seems to provide the best balance.
@@
As shown in Figure 7, adapted CLM models show less variation in performance across different masking ratios compared to fully MLM-trained models. Initial CLM pretraining appears to stabilize model weights, making adaptation more robust to masking ratio choices and yielding more consistent downstream performance less sensitive to this design parameter.
@@ 5 Continued Pretraining from CLM and MLM Models
Figure 8: Impact of performing MLM CPT on either CLM- or MLM-pretrained models (denoted as Base). CPT is conducted for 22,000 steps on 610M models with a 40% masking ratio, following 42,000 steps of initial pretraining.
Figure 9: MLM loss curves for CLM- and MLM-pretrained models across the 3 CPT compute budgets (2,000, 12,000, and 22,000 steps). Results are reported for 610M models, with MLM using a 40% masking ratio.
@@
In this section, we focus on the CPT setting and analyze whether it is more effective to adapt a CLM-pretrained model using MLM or to continue MLM training from an MLM-pretrained model.
@@
We consider a setting in which equally-sized CLM and MLM models have been pretrained on the same data. The key question is whether it is more beneficial to spend additional compute on applying MLM CPT to the CLM model or to further train the MLM model. To keep computational costs manageable, we perform a 22,000-step CPT on the 610M model using a 40% masking ratio. As shown in Figure 8, the MLM-adapted CLM model consistently achieves superior downstream performance. On TC, where CLM-only models were already strong, performance is maintained and the gap to MLM remains. For QA and IR, the gap is effectively closed, while for SC, the MLM-adapted CLM model significantly outperforms the MLM-only model.
@@
Interestingly, we observe that it is not necessary to run as many as 22,000 CPT steps to achieve comparable performance between MLM and adapted CLM. As early as 12,000 steps, the results are already strong and broadly match those of MLM-only CPT in terms of loss (Figure 9) and downstream performance (Figure 10), with better results on TC and IR, comparable performance on SC, and nearly on par on QA. In terms of trends, applying MLM CPT on a CLM model also appears more promising, showing a steeper improvement curve toward the end, while MLM-only training seems to plateau (particularly noticeable on SC).
@@ 6 Related Work
Learning universal text representations has been a central focus in NLP, with MLM emerging as the dominant pretraining objective following the success of models like BERT (Devlin et al., 2019). MLM-based encoders leverage bidirectional attention to capture rich contextual dependencies, enabling strong performance across a wide range of tasks such as classification, question answering, and named entity recognition. Subsequent work has explored architectural extensions (e.g., RoBERTa (Liu et al., 2019), DeBERTa (He et al., 2021), ModernBert (Warner et al., 2024)) and more efficient pretraining strategies (Clark et al., 2020). In contrast, decoder-only architectures such as GPT-style models were originally optimized for autoregressive generation (Brown et al., 2020; Grattafiori et al., 2024; Jiang et al., 2023; Almazrouei et al., 2023), making them less naturally suited for text representation tasks.
@@
Several recent studies have investigated the key factors influencing the performance of text representation models. BehnamGhader et al. (2024) propose a general framework for adapting decoder-only models to embedding tasks, accompanied by a series of ablations aimed at isolating the primary contributors to their performance. Some papers explore how to reintroduce bidirectional attention in causal decoder models, by repeating sequences twice (Springer et al., 2024), training with custom-shaped attention masks (Raffel et al., 2023) or selectively removing the causal attention mask on parts of the sequence in a post-training phase (Kopiczko et al., 2024; Beyer et al., 2024). Wettig et al. (2023) explore the impact of the masking ratio on downstream tasks such as sequence classification and question answering, with a focus on learning dynamics, model scale, and fine-tuning behavior. Building on this line of work, we present a comprehensive analysis of how the choice of learning objective and attention mask during training influences downstream representation quality. By employing a unified and controlled pretraining setup, we ensure fair comparisons and minimize confounding factors, allowing for more precise attribution of the effects observed.
7 Conclusion and Future Work
@@
Our study challenges the long-standing assumption that Masked Language Modeling (MLM) is the universally optimal approach for encoder pretraining. Through large-scale experiments at the 100B training token scale, we provide robust empirical evidence that encoders can be trained more data-efficiently without relying solely on MLM objectives. In particular, we demonstrate that decoder-only models trained with Causal Language Modeling (CLM) offer clear advantages in data efficiency, training stability, and performance on specific text representation tasks. Moreover, we show that sequential training (first with CLM followed by MLM) is more effective than MLM alone. These results suggest that the most effective path to high-performing encoder models may involve leveraging pretrained decoder models and continuing with MLM-based training, paving the way for future state of the art encoders at a discounted price. We hope our findings and released resources will support further advancements in this area.
Future work. Future research may extend our findings to Vision-Language Models (VLMs), many of which use decoder-based architectures, potentially improving their representation learning capabilities. Further investigations could also apply our insights to pretrain state-of-the-art encoders by capitalizing on the complementary strengths of CLM and MLM, and leveraging the availability of powerful open-source decoders, as suggested in this work.
Limitations. Our experiments are conducted at the 100B training token scale, over 5 times the compute-optimal ratio suggested for decoders in Hoffmann et al. (2022), but our loss curves suggest that further performance improvements are possible with extended training. In extremely high-compute and high-data regimes, the data-efficiency gains from CLM may diminish as they are counterbalanced by increased training capacity. Additionally, our study focuses on large-scale ablations conducted solely in English. It remains uncertain whether our conclusions generalize to other languages, where linguistic and structural differences may affect pretraining efficacy. We also do not examine the impact of common post-training practices applied to decoder models, such as fine-tuning on reasoning or mathematical datasets. Consequently, we have not analyzed how this post-training phase might influence model behavior or scaling trends relative to models pretrained from scratch with encoder-style data. Lastly, our evaluation centers on standard encoder benchmarks. Emerging use cases, such as late interaction (Khattab & Zaharia, 2020) or late chunking (Günther et al., 2024; Conti et al., 2025), which depend heavily on bidirectional attention, may surface even larger performance gaps.
Acknowledgments
We gratefully acknowledge the ADASTRA supercomputer (CINES) for providing technical support and high-performance computing (HPC) resources through projects C1615122, GDA2401, and CAD16490. We also thank the French government for its support through the France 2030 program, as part of the ArGiMi project. This work was further supported by the European Union’s Horizon Europe Research and Innovation Actions (UTTER, grant 101070631), the DECOLLAGE project (ERC-2022-CoG, grant 101088763), and the Portuguese Recovery and Resilience Plan via project C645008882-00000055 (Center for Responsible AI). Additional support was provided by FCT/MECI through national funds and, when applicable, co-funded EU funds under UID/50008: Instituto de Telecomunicações. Finally, we thank Antoine Chaffin for the insightful discussions.
@@ Appendix B Details on Evaluation Datasets
  • CoNLL (Tjong Kim Sang & De Meulder, 2003) — NER dataset comprising English and German newswire text, annotated with 4 entity types: person, organization, location, and miscellaneous. In this work, we use only the English portion.
@@ Appendix C Detailed Results
This appendix provides detailed results comparing CLM and MLM objectives, including variations in masking ratios and model sizes. Section C.1, Section C.1, Section C.1, and Section C.1 present dataset-level results for SC, TC, QA, and IR, respectively.
@@ C.4 Continued Pretraining
This appendix presents detailed results for the CPT setting, where different CPT lengths are evaluated. All results correspond to 610M-parameter models trained with a 40% masking ratio under the MLM objective. Table 14, Table 15, Table 16, and Table 17 report dataset-level performance for SC, TC, QA, and IR, respectively.

### v3

@@ 4 Two-Stage CLM+MLM Pretraining
Building on section 3, which highlights clear advantages of CLM over MLM in terms of token-level representation quality, data efficiency, and training stability, we investigate the benefits of applying CLM and MLM sequentially during pretraining. We focus on the PFS setting, where a single learning rate scheduler is used, and evaluate this approach under fixed compute budgets of 12,000, 22,000, and 42,000 training steps, using different splits: 100% CLM, 25%-75%, 50%-50%, 75%-25%, and 100% MLM. To keep computation costs manageable, we focus on the 610M model scale and use a fixed masking ratio of 40%.
Under fixed compute constraints, starting pretraining with CLM and continuing with MLM yields better results than pure MLM.
As shown in Figure 6, combining CLM and MLM consistently improves downstream performance compared to MLM-only training. In particular, the 25%-75% split reliably surpasses the MLM baseline, while even a 75% CLM allocation maintains comparable results. These findings confirm the benefits of CLM pretraining for learning strong text representations and showcase the synergies it provides when combined sequentially with MLM training.
@@
As shown in Figure 7, adapted CLM models show less variation in performance across different masking ratios compared to fully MLM-trained models. Initial CLM pretraining appears to stabilize model performance, making adaptation more robust to masking ratio choices.
@@ 5 Continued Pretraining from CLM and MLM Models
Figure 8: Impact of performing CPT on either CLM- or MLM-pretrained models (denoted as Base). CPT is conducted with MLM for 22,000 steps on 610M models with a 40% masking ratio, following 42,000 steps of initial pretraining.
@@
In this section, we focus on the CPT setting, starting from equally sized CLM and MLM models pretrained on the same data (100B tokens in total). The key question is whether it is more beneficial to allocate additional compute to applying MLM CPT to a CLM model or to further train an MLM model. We consider 3 CPT budgets: 2,000, 12,000, and 22,000 steps, and focus on the 610M model scale with a 40% masking ratio to keep computational costs manageable.
@@
As shown in Figure 8, the MLM-adapted CLM model consistently achieves superior downstream performance. On TC, where CLM-only models were already strong, performance is maintained and the gap to MLM remains. For QA and IR, the gap is effectively closed, while for SC, the MLM-adapted CLM model significantly outperforms the MLM-only model.
@@
Interestingly, we observe that it is not necessary to run as many as 22,000 CPT steps to achieve comparable performance between MLM and adapted CLM. As early as 12,000 steps, the results are already strong and broadly match those of MLM-only CPT (Figure 9), with better results on TC and IR, comparable performance on SC, and nearly on par on QA. In terms of trends, applying MLM CPT on a CLM model also appears more promising, showing a steeper improvement curve toward the end, while MLM-only training seems to plateau (particularly noticeable on SC).
@@ 6 Related Work
Learning universal text representations has been a central focus in NLP, with MLM emerging as the dominant pretraining paradigm following the success of models like BERT (Devlin et al., 2019). MLM-based encoders leverage bidirectional attention to capture rich contextual dependencies, enabling strong performance across a wide range of tasks such as classification, question answering, and named entity recognition. Subsequent work has explored architectural extensions (e.g., RoBERTa (Liu et al., 2019), DeBERTa (He et al., 2021), ModernBert (Warner et al., 2024b)) and more efficient pretraining strategies (Clark et al., 2020). In contrast, decoder-only architectures such as GPT-style models were originally optimized for autoregressive generation (Brown et al., 2020; Grattafiori et al., 2024; Jiang et al., 2023; Almazrouei et al., 2023), making them less naturally suited for text representation tasks.
@@
Several recent studies have investigated the key factors influencing the performance of text representation models. On the MLM-only side, Wettig et al. (2023) explore the impact of the masking ratio on downstream tasks such as sequence classification and question answering, with a focus on learning dynamics, model scale, and fine-tuning behavior. From another perspective, BehnamGhader et al. (2024) propose a general framework for adapting decoder-only models to embedding tasks, along with a series of ablations designed to isolate the primary contributors to performance. Complementary work explores ways to reintroduce bidirectional attention into causal decoder models, such as by repeating sequences (Springer et al., 2024), using custom-shaped attention masks (Raffel et al., 2023), or selectively removing the causal mask on portions of the sequence in a post-training phase (Kopiczko et al., 2024; Beyer et al., 2024). Articles most closely related to ours propose controlled setups for comparing CLM, MLM, and hybrid strategies, but have either focused on very small amounts of pretraining data (Charpentier and Samuel, 2024), or prioritized other design parameters. For example, Weller et al. (2025) train on very large data regimes (2T tokens), analyze MLM-then-CLM learning curricula, and evaluate on generative benchmarks, providing practical insights into the impact of training design on downstream performance. However, their approach does not support extensive evaluation on encoder-specific tasks and limits exploration of additional training design choices, such as masking ratio or the CLM-MLM share in pretraining. Building on these efforts, we provide a comprehensive analysis of how the choice of learning objective and attention mask during training affects downstream representation quality. By employing a unified and controlled pretraining setup and evaluating across a broad range of representation tasks with statistically grounded results, we ensure fair comparisons and reduce confounding factors, enabling more precise attribution of observed effects.
7 Conclusion
@@
Our study challenges the long-standing assumption that Masked Language Modeling (MLM) is the universally optimal approach for encoder pretraining. Through large-scale experiments at the 100B training token scale, we provide robust empirical evidence that encoders can be trained more data-efficiently without relying solely on MLM. In particular, we demonstrate that decoder-only models trained with Causal Language Modeling (CLM) offer clear advantages in data efficiency, training stability, and performance on specific text representation tasks. Moreover, we show that sequential training (first with CLM followed by MLM) is more effective than MLM alone. These results suggest that the most effective path to high-performing encoder models may involve leveraging pretrained decoder models and continuing with MLM-based training, paving the way for future state of the art encoders at a discounted price. We hope our findings and released resources will support further advancements in this area.
8 Limitations and Future Work
Our study deliberately focuses on a targeted set of parameters, namely learning objective, model size, training setup, data budget, and masking ratio. As in large-scale studies such as (Kaplan et al., 2020; Hoffmann et al., 2022), this required fixing several other parameters, including the model architecture, tokenizer, language, and training mixture. While these factors are certainly of interest, varying them would introduce significant computational overhead, making them difficult to incorporate into our current scope. We therefore view them as a promising direction for future work, including extensions to multilingual settings or to additional modalities such as vision–language.
Additionally, although we provide insights into scaling behaviors with respect to both model size and data budget, we constrained these dimensions to 1B parameters and 100B training tokens. This allowed us to keep compute costs manageable while still exploring a broad range of design parameters. While 100B tokens correspond to roughly five times the compute-optimal budget suggested by Hoffmann et al. (2022) for 1B-parameter decoder models, and most pre-trained encoders in the literature remain below the 1B scale (Devlin et al., 2019; Liu et al., 2019; He et al., 2021; Warner et al., 2024a), we still view pushing beyond these limits as an important direction for future work, as many top entries in the MTEB benchmark (Muennighoff et al., 2023; Enevoldsen et al., 2025) frequently exceed the 1B-parameter range (Zhang et al., 2025; Babakhin et al., 2025). Exploring more sophisticated training curricula also appears promising for improving downstream performance, for instance, alternating CLM and MLM objectives multiple times throughout pretraining.
In a similar spirit, our study focuses on the extensive evaluation of pre-trained models and deliberately does not include contrastive post-training for zero-shot retrieval. Applying such post-training across a large number of training configurations would have incurred substantial computational costs and could have introduced additional confounding factors that require careful control (Xiong et al., 2020; Qu et al., 2021; Wang et al., 2022; Mohr et al., 2024; Chen et al., 2024; Sturua et al., 2024). While this lies outside the scope of the present work, exploring contrastive post-training would be a valuable way to complement and extend our findings.
Finally, several observations emerged throughout the paper that appear intriguing but fall outside the current scope of our study. Future work could, for example, further investigate the relationship between masking ratio and downstream performance across tasks. This includes providing deeper insights into patterns such as the U-shaped curve observed in Figure 3 for TC on the 610M and 1B models, in contrast with the monotonic trend observed for IR. More broadly, gaining a deeper understanding of how pretraining learning objectives transfer to different downstream tasks remains an interesting yet largely unexplored direction with substantial promise.
Ethics Statement
Environmental and compute considerations.
This work examines different pretraining strategies for encoder models, with the aim of exploring approaches that may reduce computational costs while still producing useful text representations. By considering alternatives to traditional Masked Language Modeling (MLM) and investigating a biphasic pretraining strategy, we seek ways to train encoder models in a more data- and compute-efficient manner. Reducing the computational resources required for model training can have both environmental and practical benefits. More efficient encoders may help lower energy consumption and the associated footprint of large-scale NLP training. Additionally, effective encoder models could in some cases reduce the need to rely on large generative models for downstream tasks, thereby decreasing dependence on more compute-intensive systems.
Responsible use of LLMs.
In preparing this manuscript, we occasionally used suggestions from LLMs (GPT-5) to guide improvements in clarity, grammar, and overall readability. All scientific content, including experimental design, codebase, data analysis, results, and interpretations, is independently developed by the authors. LLMs are not involved in generating, modifying, or interpreting any experimental results, nor in producing code or analyses. Their use is strictly limited to selectively refining language to ensure clear and effective communication of our research.
Reproducibility Statement
We have made every effort to ensure that our experiments are reproducible. All pretraining and evaluation experiments are described in detail in the paper, including model architectures, training objectives, data sources, and hyperparameter settings. To facilitate replication, we release all project artifacts, including pretrained models, training scripts, and evaluation code.
@@ Appendix B Details on Evaluation Datasets
  • CoNLL (Tjong Kim Sang and De Meulder, 2003) — NER dataset comprising English and German newswire text, annotated with 4 entity types: person, organization, location, and miscellaneous. In this work, we use only the English portion.
@@ Appendix C Detailed Results
This appendix provides detailed results comparing CLM and MLM pretraining, including variations in masking ratios and model sizes. subsection C.1, subsection C.1, subsection C.1, and subsection C.1 present dataset-level results for SC, TC, QA, and IR, respectively.
@@ C.4 Continued Pretraining
This appendix presents detailed findings for the CPT setting, evaluating different CPT lengths. All experiments use 610M-parameter models, with MLM reported for a 40% masking ratio. Table 14, Table 15, Table 16, and Table 17 summarize dataset-level performance for SC, TC, QA, and IR, respectively.
@@
Appendix D Additional Results
This appendix provides additional results that complement our main analysis. Figure 10 complements Figure 4 by showing that CLM is not only more data-efficient than MLM but also more efficient in terms of training FLOPs. Figure 11 confirms, at the 1B scale, the same trends observed for the 610M model in Figure 6. Figure 12 presents loss curves across different CPT training budgets, starting from either an MLM- or CLM-pre-trained model and then continuing with MLM.
Figure 10: Downstream performance with respect to pretraining FLOPs for CLM and MLM. Results are reported for 610M models, with a 40% masking ratio for MLM. M. Training FLOPs are computed following the formula provided by Hoffmann et al. (2022).
Figure 11: Impact of two-stage CLM+MLM pretraining on downstream performance after 42,000 training steps. Results are reported for the 1B model, with MLM applied using a 40% masking ratio. To keep computation costs manageable, CLM+MLM results are shown only for the 25%–75% regime.
Figure 12: Loss trends in the CPT setting. The blue curve shows the loss trajectory of a model using MLM for both pretraining and continued pretraining, while the red curve depicts a CLM model adapted using the MLM objective during CPT. Results are reported for the 610M model with a 40% masking ratio for MLM.